Consider the following ordinary differential equation: y - sin(4t) = 0 (Eq. 4) The boundary condition is that y(0) = -0.25.
Answer #1

Taylor Method to solve ODE: MATLAB CODE clear all symst yO=-0.25 h=0.1; dy-sin(4*t); d2y=diff (dy,t); d3y=diff (d2y,t); d4y=d